Are circular coil of radius 40mm is found to produce a field 12.1×10^-6 T at a point 30mm along its axis.(a) what is the electromagnetic moment? (b) what maximum torque could it experience in a uniform magnetic field of 3.0×10^-3T.


1
Expert's answer
2021-02-11T17:11:00-0500

a) Electromagnetic moment of the coil is given as:


"m = IS"

where "S = \pi R^2" is the area of the coil, where "R = 40mm = 0.04m" is its radius, and "I" is the current in the coil. In turn the current can be found from the following expression:


"B = \dfrac{\mu_0I}{2R}\left( 1+\dfrac{h^2}{R^2} \right)^{-3/2}"

where "B = 12.1\times 10^{-6}T" is the bagnetic field, "h =30mm = 0.03m" is the distance along the axis, and "\mu_0 =4\pi\times 10^{-7}H/m" is the magnetic permeability. Expressing the current, obtain:


"I = \dfrac{2RB}{\mu_0}\left( 1+\dfrac{h^2}{R^2} \right)^{3/2}"

And the moment:


"m = \dfrac{2\pi R^3B}{\mu_0}\left( 1+\dfrac{h^2}{R^2} \right)^{3/2}\\
m = \dfrac{2\pi \cdot 0.04^3\cdot 12.1\times 10^{-6}}{4\pi\cdot 10^{-7}}\left( 1+\dfrac{0.03^2}{0.04^2} \right)^{3/2} \approx 0.0076\space A\cdot m^2"

b) The maximum torque is given as follows:


"\tau = mB_0"

where "B_0 = 3\times 10^{-3}T" is the external field. Thus, obtain:


"\tau \approx 2.3\times 10^{-5}\space N\cdot m"

Answer. a) "0.0076\space A\cdot m^2" b) "2.3\times 10^{-5}\space N\cdot m".